OpenStudy (anonymous):

Help? I need to find the domain, range, x and y intercepts, and the horizontal and vertical asymptotes for both of these functions.

OpenStudy (anonymous):

What are the functions?

OpenStudy (anonymous):

\[1) \frac{ 2 }{ x^2-2x-3 }\]

OpenStudy (anonymous):

\[2) \frac{ x^2+x-2 }{ x^2-3x-4}\]

OpenStudy (anonymous):

domain all real numbers

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

Really? The first one looks weird on my calculator.

OpenStudy (anonymous):

nvmd, 1st one is not x=4 or x=-1

OpenStudy (anonymous):

range is y<=8/25 and y>0

OpenStudy (anonymous):

-8/25***

OpenStudy (anonymous):

Wait, are you saying the first one is x=4 and x=-1, or it is not equal to x=4, an -1?

OpenStudy (anonymous):

all real numbers except those 2

OpenStudy (anonymous):

no x int and y int is -1/2

OpenStudy (anonymous):

hor asym at +/_ infinity vert asym at -1 and 4

OpenStudy (anonymous):

haha i screwed that all up. i took the numerator of your first equation and denominator of your second

OpenStudy (anonymous):

First equation- Domain is all real numbers except -1 and 3, range is y<=-1/2 and y>0, no x int, y int=-2/3, hor asymptotes at +/- infinity and vert at -1 and 3

OpenStudy (anonymous):

Ah, that makes more sense. :)